# Regularity Propagation of Global Weak Solutions to a Navier–Stokes–Cahn–Hilliard System for Incompressible Two-phase Flows with Chemotaxis and Active Transport

### v1

@@ 2.3 Proof Strategies
We conclude by discussing the applied techniques and strategies of proof.
2.3.1 Existence of global weak solutions in $[0,+\infty)$
Regarding Proposition 2.1, for an arbitrary but fixed final time $T\in(0,+\infty)$, the existence of a weak solution $(\bm{v},\varphi,\mu,\sigma)$ to problem (1.1)–(1.3) on $[0,T]$ can be established through a suitable semi-Galerkin scheme and a compactness argument akin to [27, Section 3], with minor modifications. However, owing to the lack of uniqueness for weak solutions in three dimensions, in order to obtain the global existence on the unbounded interval $[0,+\infty)$, one cannot simply extend the solution defined on $[0,T]$ by letting $T\to+\infty$. Instead, we observe that the approximate solutions $(\bm{v}^{m},\varphi^{m},\mu^{m},\sigma^{m})$ constructed via the semi-Galerkin scheme as in [27] are indeed unique ($m\in\mathbb{Z}^{+}$ being the parameter in the Galerkin approximation). Furthermore, these solutions satisfy the mass relation and energy equality on the corresponding existence interval $[0,T_{m}]$:
@@ 2.3.1 Existence of global weak solutions in $[0,+\infty)$
and $\mathcal{F}$ is defined as in (1.8). Based on (2.5) and (2.6), for every fixed $m\in\mathbb{Z}^{+}$, we can extend the approximate solution $(\bm{v}^{m},\varphi^{m},\mu^{m},\sigma^{m})$ from the finite interval $[0,T_{m}]$ to $[0,+\infty)$. Since (2.5) and (2.6) provide sufficient estimates for $(\bm{v}^{m},\varphi^{m},\mu^{m},\sigma^{m})$ that are uniform with respect to $m$ (see also the lower-order energy estimates in [30, Section 3.2]), we can employ a compactness argument and a diagonal extraction process analogous to that in [8, Chapter V, Section 1.3.6] for the three-dimensional Navier–Stokes equations. This allows us to find a subsequence of $\big{\{}(\bm{v}^{m},\varphi^{m},\mu^{m},\sigma^{m})\big{\}}$ that converges, as $m\to+\infty$, to a weak solution of problem (1.1)–(1.3) on the interval $[0,N]$ for every $N\in\mathbb{Z}^{+}$. Consequently, the limit of this convergent subsequence yields a global weak solution to problem (1.1)–(1.3) defined on $[0,+\infty)$, possessing the required regularity properties stated in Proposition 2.1. A similar approach will be utilized in the proof of Theorem 3.1 below.
From the mass relation (2.5) (i.e., $\overline{\varphi^{m}}$ is bounded and indeed independent of $m$) and the fact that for any $T\in(0,+\infty)$ the approximate chemical potential $\mu^{m}$ weakly converges (up to a subsequence) in $L^{2}(0,T;L^{2}(\Omega))$, we find that for all $0\leq t_{1}<t_{2}\leq T$, it holds
@@
as $m\to+\infty$ (again up to a subsequence). This observation, combined with the energy equality (2.6) for the approximate solutions $(\bm{v}^{m},\varphi^{m},\mu^{m},\sigma^{m})$, and a standard argument for the Navier–Stokes–Cahn–Hilliard system (see, e.g., [1, Theorem 1]) enables us to conclude the energy inequality (2.3). Therefore, the details of the proof for Proposition 2.1 are omitted in this paper.
@@ 2.3.2 Global regularity
The proof of Theorem 2.1 is inspired by the recent work [2]. One key point in that work is to establish full regularity for the solution to the Cahn–Hilliard equation with the logarithmic potential (1.5) and a divergence-free drift, under the weak assumption $\bm{v}\in L^{2}(0,+\infty;\bm{H}^{1}_{0,\mathrm{div}}(\Omega))$ (see [2, Theorem 2.4]). For our problem (1.1)–(1.3), to address the coupling due to chemotaxis and mass transport, we need to investigate the regularity of weak solutions to a convective Cahn–Hilliard–diffusion system for $(\varphi,\sigma)$ with a prescribed divergence-free velocity $\bm{v}\in L^{2}(0,+\infty;\bm{H}^{1}_{0,\mathrm{div}}(\Omega))$, as seen in (3.1)–(3.3). Detailed analysis will be presented in Section 3. Here, new challenges emerge in the equation for $\sigma$ due to the low regularity of $\bm{v}$ and the cross diffusion structure. This leads to limited regularity of $\sigma$, which subsequently hinders the attainment of higher-order spatial regularity for $\varphi$ through (3.1b). We first prove the existence of global weak solutions to problem (3.1)–(3.3) along with a conditional uniqueness result, as outlined in Theorem 3.1. Subsequently, in Theorem 3.2, we establish the existence and uniqueness of more regular solutions under stronger assumptions on the initial data, with $\bm{v}$ still belonging to the Leray–Hopf class. These results imply the instantaneous regularization of weak solutions $(\varphi,\mu,\sigma)$ after any arbitrary positive time. Although we can recover the regularity of $\varphi$ as in [2], there is a loss of regularity for $\sigma$ (due to the low regularity of $\bm{v}$). For instance, with the initial datum $\sigma_{0}\in H^{1}(\Omega)$, we only obtain $\sigma\in L^{\infty}(0,+\infty,L^{6}(\Omega))$ instead of $\sigma\in L^{\infty}(0,+\infty,H^{1}(\Omega))$.
@@
Next, an essential step towards achieving the eventual regularity of $\bm{v}$ is to investigate the long-time behavior of global weak solutions through the $\omega$-limit set, see Proposition 4.1. Specifically, we demonstrate the strict separation of $\varphi$ from the pure states ±$\pm 1$ after a sufficiently large time. This property is crucial, as it ensures that the singular potential $\varPsi$ and its derivatives become smooth and bounded afterwards. This further enables us to derive a weak-strong uniqueness result for the full system (1.1)–(1.3) by the relative entropy method, see Proposition 4.2. By combining the result on local strong well-posedness obtained in [30], we find that every weak solution transitions into a strong one for sufficiently large times. Subsequently, global regularity of $\bm{v}$ is achieved by applying the theory on global well-posedness for the Navier–Stokes system with variable viscosity [1]. Finally, with the improved regularity of $\bm{v}$, we can show that the chemical density $\sigma$ will regularizes again, in turn enhances the spatial regularity of $\varphi$.
@@
2.3.3 Convergence to equilibrium
Theorem 2.1 implies that every global weak solution $(\bm{v},\varphi,\mu,\sigma)$ to problem (1.1)–(1.3) transitions into a global strong one after a sufficiently large time, with uniform-in-time estimates in $\bm{H}^{1}_{0,\mathrm{div}}(\Omega)\times H^{3}(\Omega)\times H^{1}(\Omega)\times H^{1}(\Omega)$. Furthermore, under the additional assumption that $\varPsi$ is real analytic in $(-1,1)$, we can apply an extended Łojasiewicz–Simon type inequality (see [30, Theorem 4.1]) to demonstrate that $(\bm{v},\varphi,\sigma)$ converges to a single equilibrium $(\mathbf{0},\varphi_{\infty},\sigma_{\infty})$ as $t\to+\infty$. The proof of Theorem 2.2 follows the same argument as that for [30, Theorem 2.2], thus its details are omitted in this paper.
@@ 3 A Cahn–Hilliard–Diffusion System with Divergence-free Drift
In (3.1), the prescribed divergence-free velocity field $\bm{v}$ is suppose to satisfy the minimum regularity, that is, $\bm{v}\in L^{2}(0,+\infty;\bm{H}_{0,\mathrm{div}}^{1}(\Omega))$.
The first theorem gives the existence and conditional uniqueness of global weak solutions:
@@
Let $\Omega$ be a bounded domain in $\mathbb{R}^{3}$, with boundary $\partial\Omega$ of class $C^{2}$. Suppose that the hypotheses (H2)–(H3) are satisfied. In addition, we assume
@@
    $\varphi\in L^{\infty}(0,+\infty;H^{1}(\Omega))\cap L^{4}_{\mathrm{uloc}}(0,+\infty;H^{2}_{N}(\Omega))\cap L^{2}_{\mathrm{uloc}}(0,+\infty;W^{2,6}(\Omega)),$
    $\partial_{t}\varphi\in L^{2}(0,+\infty;(H^{1}(\Omega))^{\prime}),\quad\varPsi^{\prime}(\varphi)\in L^{2}_{\mathrm{uloc}}(0,+\infty;L^{6}(\Omega))$
@@
    $\mu\in L_{\mathrm{uloc}}^{2}(0,+\infty;H^{1}(\Omega)),\quad\nabla\mu\in L^{2}(0,+\infty;\bm{L}^{2}(\Omega)),$
    $\sigma\in L^{\infty}(0,+\infty;L^{2}(\Omega))\cap L_{\mathrm{uloc}}^{2}(0,+\infty;H^{1}(\Omega))\cap W^{1,\frac{4}{3}}_{\mathrm{uloc}}(0,+\infty;(H^{1}(\Omega))^{\prime}).$
@@
    $\left\langle\partial_{t}\varphi,\xi\right\rangle_{(H^{1})^{\prime},H^{1}}+({\bm{v}\cdot\nabla\varphi},\xi)=-(\nabla\mu,\nabla\xi)-\alpha(\varphi-c_{0},\xi),$ (3.4a)
    $\left\langle\partial_{t}\sigma,\xi\right\rangle_{(H^{1})^{\prime},H^{1}}+(\bm{v}\cdot\nabla\sigma,\xi)+(\nabla\sigma,\nabla\xi)=\chi(\nabla\varphi,\nabla\xi),$ (3.4b)
@@
    $\varphi|_{t=0}=\varphi_{0},\quad\sigma|_{t=0}=\sigma_{0},\quad\text{a.e. in}\ \Omega.$
@@
    Conditional uniqueness. For any $T\in(0,+\infty)$, let $(\varphi_{i},\mu_{i},\sigma_{i})$, $i=1,2$, be two weak solutions to problem (3.1)–(3.3) on $[0,T]$ corresponding to the same initial datum $(\varphi_{0},\sigma_{0})$ as in (1). If in addition,
    $\sigma_{2}\in L^{8}(0,T;L^{4}(\Omega)),$
    then $(\varphi_{1}(t),\mu_{1}(t),\sigma_{1}(t))=(\varphi_{2}(t),\mu_{2}(t),\sigma_{2}(t))$ for all $t\in[0,T]$.
Next, we show that under additional assumptions on the domain and initial data, the solution becomes more regular and unique.
@@
Let $\Omega$ be a bounded domain in $\mathbb{R}^{3}$, with boundary $\partial\Omega$ of class $C^{3}$. Assume that the hypotheses (H2)–(H3) are satisfied and
$\bm{v}\in L^{2}(0,+\infty;\bm{H}_{0,\mathrm{div}}^{1}(\Omega)).$
@@
$\varphi\in L^{\infty}(0,+\infty;W^{2,6}(\Omega)),\qquad\partial_{t}\varphi\in L^{2}(0,+\infty;H^{1}(\Omega)),$
@@
$\mu\in L^{\infty}(0,+\infty;H^{1}(\Omega))\cap L_{\mathrm{uloc}}^{2}(0,+\infty;H^{3}(\Omega)),$
$\sigma\in L^{\infty}(0,+\infty;L^{6}(\Omega))\cap L_{\mathrm{uloc}}^{2}(0,+\infty;H^{1}(\Omega))\cap H^{1}(0,+\infty;(H^{1}(\Omega))^{\prime}).$
@@
for almost all $t\in(0,+\infty)$ and all $\xi\in H^{1}(\Omega)$. The boundary condition (3.2) holds almost everywhere on $\partial\Omega\times(0,+\infty)$ and the initial condition (3.3) is fulfilled almost everywhere in $\Omega$. Besides, the solution $(\varphi,\mu,\sigma)$ satisfies the energy equality
@@
$\quad=\mathcal{F}(\varphi(t_{1}),\sigma(t_{1}))-\alpha\int_{t_{1}}^{t_{2}}\!\int_{\Omega}(\overline{\varphi}-c_{0})\mu\,\mathrm{d}x\mathrm{d}s+\int_{t_{1}}^{t_{2}}\!\int_{\Omega}\bm{v}\cdot\nabla(\mu+\chi\sigma)\varphi\,\mathrm{d}x\mathrm{d}s,$ (3.8)
for all $0\leq t_{1}<t_{2}<+\infty$, where $\mathcal{F}(\varphi,\sigma)$ is given by (1.8). If in addition, $\bm{v}\in L^{\infty}(0,+\infty;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))$, then we have $\partial_{t}\varphi\in L^{\infty}(0,+\infty;(H^{1}(\Omega))^{\prime})$.
@@
For the single Cahn–Hilliard equation with non-smooth potential and divergence-free drift, existence and uniqueness of global weak solutions were shown in [1, Theorem 6] and regularity under minimum requirement on the velocity was established in [2, Theorem 2.4]. Theorem 3.1 and Theorem 3.2 generalize those previous results to the Cahn-Hilliard-diffusion system (3.1) with cross diffusion structure as well as nonlocal interaction of Oono’s type. Under the condition $\bm{v}\in L^{2}(0,+\infty;\bm{H}_{0,\mathrm{div}}^{1}(\Omega))$ we can still gain full regularity of $(\varphi,\mu)$, but there is a loss in the regularity of $\sigma$.

### v2

@@ 2.3 Proof strategies
We conclude this section by discussing the applied techniques and proof strategies.
2.3.1 Existence of global weak solutions in $0,+\infty)$
Regarding Proposition 2.1, for an arbitrary but fixed final time $T\in(0,+\infty)$, the existence of a weak solution $(\bm{v},\varphi,\mu,\sigma)$ to problem (1.1)–(1.3) on $[0,T]$ can be established through a suitable semi-Galerkin scheme and a compactness argument similar to [31, Section 3], with minor modifications. However, owing to the lack of uniqueness for weak solutions in three dimensions, in order to obtain the global existence on the unbounded interval $0,+\infty)$, one cannot simply extend the solution defined on $[0,T]$ by letting $T\to+\infty$. Instead, we observe that the approximate solutions $(\bm{v}^{m},\varphi^{m},\mu^{m},\sigma^{m})$ constructed via the semi-Galerkin scheme and Schauder’s fixed-point theorem as in [31] are indeed unique ($m\in\mathbb{Z}^{+}$ being the parameter in the Galerkin approximation). Furthermore, these solutions satisfy the following mass relations and the energy equality on their corresponding existence interval $0,T_{m}]$:
@@ 2.3.1 Existence of global weak solutions in $0,+\infty)$
and $\mathcal{F}$ is defined as in (1.8). If $\alpha=0$, the total energy decreases with respect to time. If $\alpha>0$, applying the same argument for [34, (3.20)], we can control the reminder term on the right-hand side of (2.8) by
$-\alpha\int_{\Omega}(\overline{\varphi^{m}}-c_{0})\mu^{m}\,\mathrm{d}x\leq\frac{1}{2}\big(\|\nabla\mu^{m}\|^{2}+\|\nabla(\sigma^{m}-\chi\varphi^{m})\|^{2}\big)+C(1+\alpha^{3})\alpha e^{-\alpha t}|\overline{\varphi_{0}}-c_{0}|,$
for almost all $t\in[0,T_{m}]$, where $C>0$ depends on $\overline{\varphi_{0}}$, $\overline{\sigma_{0}}$, $\Omega$ and the coefficients of the system except $\alpha$. We note that the first two terms on the right-hand side of the above inequality can be absorbed into the left-hand side of (2.8) (after integration on $t_{1},t_{2}]$), while the last term belongs to $L^{1}(0,+\infty)$. Thus, based on (2.7) and (2.8), for every fixed $m\in\mathbb{Z}^{+}$, we can obtain uniform estimates with respect to time for the approximate solution $(\bm{v}^{m},\varphi^{m},\mu^{m},\sigma^{m})$ and extend it from the finite interval $0,T_{m}]$ to $0,+\infty)$. Similar arguments can be found in the proof of Theorem 3.1 below. On the other hand, since (2.7) and (2.8) provide sufficient estimates for $(\bm{v}^{m},\varphi^{m},\mu^{m},\sigma^{m})$ that are uniform with respect to the approximating parameter $m$ (see also the lower-order energy estimates in [34, Section 3.2]), we can employ a compactness argument and a diagonal extraction process analogous to that in [10, Chapter V, Section 1.3.6] for the three-dimensional Navier–Stokes equations. This allows us to find a subsequence of $\big\{(\bm{v}^{m},\varphi^{m},\mu^{m},\sigma^{m})\big\}$ that converges, as $m\to+\infty$, to a weak solution of problem (1.1)–(1.3) on the interval $[0,N]$ for every $N\in\mathbb{Z}^{+}$. Consequently, the limit of this convergent subsequence yields a global weak solution to problem (1.1)–(1.3) defined on $0,+\infty)$, possessing the required regularity properties stated in Proposition 2.1.
From the mass relation (2.7) (i.e., $\overline{\varphi^{m}}$ is bounded and indeed independent of $m$) and the fact that for any $T\in(0,+\infty)$ the approximate chemical potential $\mu^{m}$ weakly converges (up to a subsequence) in $L^{2}(0,T;L^{2}(\Omega))$, we find that for all $0\leq t_{1}\leq t_{2}\leq T$, it holds
@@
as $m\to+\infty$ (again up to a subsequence). This observation, combined with the energy equality (2.8) for the approximate solutions $(\bm{v}^{m},\varphi^{m},\mu^{m},\sigma^{m})$, and a standard argument for the Navier–Stokes–Cahn–Hilliard system (see, e.g., [1, Theorem 1]) enables us to conclude the energy inequality (2.5).
Taking into account the above discussions, the details of the proof for Proposition 2.1 are omitted in this paper.
@@ 2.3.2 Global regularity
The proof of Theorem 2.1 is inspired by the recent work [3]. One key point therein is to establish full regularity for the solution of the Cahn–Hilliard equation with the logarithmic potential (1.5) and a divergence-free drift, under the weak assumption $\bm{v}\in L^{2}(0,+\infty;\bm{H}^{1}_{0,\mathrm{div}}(\Omega))$ (see [3, Theorem 2.4]). For our problem (1.1)–(1.3), in order to address the coupling due to chemotaxis and mass transport, we need to investigate the regularity of weak solutions to a convective Cahn–Hilliard–diffusion system for $(\varphi,\mu,\sigma)$ with a prescribed divergence-free velocity $\bm{v}\in L^{2}(0,+\infty;\bm{H}^{1}_{0,\mathrm{div}}(\Omega))$, as seen in (3.1)–(3.3). Detailed analysis will be presented in Section 3. Here, new challenges emerge in the second-order parabolic equation for $\sigma$ due to the low regularity of $\bm{v}$ and the cross-diffusion structure. This leads to a limited regularity of $\sigma$, which subsequently hinders the attainment of a higher-order spatial regularity for $\varphi$ through (3.1b). We first prove the existence of global weak solutions to problem (3.1)–(3.3) along with a conditional uniqueness result, as outlined in Theorem 3.1. Subsequently, in Theorem 3.2, we establish the existence and uniqueness of more regular solutions under stronger assumptions on the initial data, with $\bm{v}$ belonging to the Leray–Hopf class, i.e.,
$\bm{v}\in L^{\infty}(0,+\infty;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))\cap L^{2}(0,+\infty;\bm{H}^{1}_{0,\mathrm{div}}(\Omega)),$
which coincides with the regularity of the global weak solution to the hydrodynamic problem (1.1)–(1.3). To prove Theorem 3.2, we study two decoupled auxiliary systems for $\sigma$ and $(\varphi,\mu)$, respectively, and then apply a bootstrap argument. These results imply the instantaneous regularization of global weak solutions $(\varphi,\mu,\sigma)$ to problem (1.1)–(1.3) after an arbitrary positive time. Although we can recover the full regularity of $\varphi$ as in [3], there is a loss of regularity for $\sigma$ (due to the low regularity of $\bm{v}$). For example, with the initial datum $\sigma_{0}\in H^{1}(\Omega)$, we only obtain $\sigma\in L^{\infty}(0,+\infty,L^{6}(\Omega))$ instead of $\sigma\in L^{\infty}(0,+\infty,H^{1}(\Omega))$.
@@
Next, an essential step towards achieving the eventual regularity of $\bm{v}$ is to investigate the long-time behavior of global weak solutions through the $\omega$-limit set, see Proposition 4.1. Specifically, we demonstrate the strict separation of $\varphi$ from the pure states $\pm 1$ after a sufficiently large time. The eventual separation property is crucial because it ensures that the singular potential $\varPsi$ and its derivatives become smooth and bounded afterwards. This property further enables us to derive a weak-strong uniqueness result for the full system (1.1)–(1.3) via the relative energy method, see Proposition 4.2. By combining the result on local strong well-posedness obtained in [34], we find that every weak solution transitions into a strong one for sufficiently large time. Subsequently, the global regularity of $\bm{v}$ is achieved by applying the theory of global well-posedness for the Navier–Stokes system with variable viscosity [1]. At last, with the improved regularity of $\bm{v}$, we can show that the chemical density $\sigma$ will regularize again, in turn improving the spatial regularity of $\varphi$.
@@
2.3.3 Convergence to a single equilibrium
Theorem 2.1 implies that every global weak solution $(\bm{v},\varphi,\mu,\sigma)$ to problem (1.1)–(1.3) becomes a global strong one after a sufficiently large time, with uniform-in-time estimates in $\bm{H}^{1}_{0,\mathrm{div}}(\Omega)\times H^{3}(\Omega)\times H^{1}(\Omega)\times H^{1}(\Omega)$. Furthermore, under the additional assumption that $\varPsi$ is real analytic in (1,1)(-1,1), we can apply an extended Łojasiewicz–Simon type inequality (see [34, Theorem 4.1]) to demonstrate that the solution $(\bm{v},\varphi,\sigma)$ converges to a single equilibrium $(\mathbf{0},\varphi_{\infty},\sigma_{\infty})$ as $t\to+\infty$. The proof of Theorem 2.2 follows the same argument as that for [34, Theorem 2.2], so its details are omitted in this paper.
@@ 3 A Cahn–Hilliard–Diffusion System with Divergence-free Drift
In (3.1), the prescribed divergence-free velocity field $\bm{v}$ is supposed to satisfy a mild assumption on its regularity, that is,
$\bm{v}\in L^{\infty}(0,+\infty;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))\cap L^{2}(0,+\infty;\bm{H}_{0,\mathrm{div}}^{1}(\Omega)).$
The first theorem establishes the existence and conditional uniqueness of global weak solutions to problem (3.1)–(3.3).
@@
Let $\Omega$ be a bounded domain in $\mathbb{R}^{3}$, with boundary $\partial\Omega$ of class $C^{2}$. Suppose that hypotheses (H2)–(H3) are satisfied and
@@
    $\varphi\in L^{\infty}(0,+\infty;H^{1}(\Omega))\cap L^{4}_{\mathrm{uloc}}([0,+\infty);H^{2}_{N}(\Omega))\cap L^{2}_{\mathrm{uloc}}([0,+\infty);W^{2,6}(\Omega)),$
    $\partial_{t}\varphi\in L^{2}(0,+\infty;(H^{1}(\Omega))^{\prime}),\quad\varPsi^{\prime}(\varphi)\in L^{2}_{\mathrm{uloc}}([0,+\infty);L^{6}(\Omega))$
@@
    $\mu\in L_{\mathrm{uloc}}^{2}([0,+\infty);H^{1}(\Omega)),\quad\nabla\mu\in L^{2}(0,+\infty;\bm{L}^{2}(\Omega)),$
    $\sigma\in L^{\infty}(0,+\infty;L^{2}(\Omega))\cap L_{\mathrm{uloc}}^{2}([0,+\infty);H^{1}(\Omega))\cap W^{1,\frac{4}{3}}_{\mathrm{uloc}}([0,+\infty);(H^{1}(\Omega))^{\prime}),$
    $\nabla(\sigma-\chi\varphi)\in L^{2}(0,+\infty;\bm{L}^{2}(\Omega)).$
@@
    $\left\langle\partial_{t}\varphi,\xi\right\rangle_{(H^{1})^{\prime},H^{1}}+({\bm{v}\cdot\nabla\varphi},\xi)=-(\nabla\mu,\nabla\xi)-\alpha(\overline{\varphi}-c_{0},\xi),$
    $\left\langle\partial_{t}\sigma,\xi\right\rangle_{(H^{1})^{\prime},H^{1}}+(\bm{v}\cdot\nabla\sigma,\xi)+(\nabla\sigma,\nabla\xi)=\chi(\nabla\varphi,\nabla\xi),$
@@
    $\mu=-\Delta\varphi+\varPsi^{\prime}(\varphi)-\chi\sigma+\beta\mathcal{N}(\varphi-\overline{\varphi}),\quad\text{a.e. in}\ \Omega\times(0,+\infty).$
@@
    $\varphi|_{t=0}=\varphi_{0},\quad\sigma|_{t=0}=\sigma_{0},\quad\text{a.e. in}\ \Omega,$
    and the solution $(\varphi,\sigma)$ satisfies the following energy inequality:
    $\frac{1}{2}\|\sigma(t)\|^{2}+\int_{0}^{t}\|\nabla\sigma\|^{2}\,\mathrm{d}s\leq\frac{1}{2}\|\sigma_{0}\|^{2}+\chi\int_{0}^{t}\!\int_{\Omega}\nabla\sigma\cdot\nabla\varphi\,\mathrm{d}x\mathrm{d}s,\quad\forall\,t>0.$ (3.5)
@@
    Conditional uniqueness. For any $T\in(0,+\infty)$, let $(\varphi_{i},\mu_{i},\sigma_{i})$, $i=1,2$, be two weak solutions to problem (3.1)–(3.3) on $[0,T]$ corresponding to the same initial datum $(\varphi_{0},\sigma_{0})$ as described in (1). Assume, in addition, $\bm{v}\in L^{\infty}(0,+\infty;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))$ and $\sigma_{2}\in L^{8}(0,T;L^{4}(\Omega))$, then
    $(\varphi_{1}(t),\mu_{1}(t),\sigma_{1}(t))=(\varphi_{2}(t),\mu_{2}(t),\sigma_{2}(t)),\quad\forall\,t\in[0,T].$
Remark 3.1.
Similarly to Remark 2.2, from the regularity of $(\varphi,\sigma)$ and using [6, Theorem 4.10.2, Chapter III], [2, Lemma 4.1], we can conclude $\varphi\in BC_{w}([0,+\infty);H^{1}(\Omega))$ and $\sigma\in BC_{w}([0,+\infty);L^{2}(\Omega))$, so that the initial data can be achieved.
Next, we show that, under additional assumptions on the domain and initial data, the weak solution becomes more regular and unique.
@@
Let $\Omega$ be a bounded domain in $\mathbb{R}^{3}$, with boundary $\partial\Omega$ of class $C^{3}$. Suppose that hypotheses (H2)–(H3) are satisfied and
$\bm{v}\in L^{\infty}(0,+\infty;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))\cap L^{2}(0,+\infty;\bm{H}_{0,\mathrm{div}}^{1}(\Omega)).$
@@
$\varphi\in L^{\infty}(0,+\infty;W^{2,6}(\Omega)\cap H^{2}_{N}(\Omega)),$
$\partial_{t}\varphi\in L^{2}_{\mathrm{uloc}}(0,+\infty;H^{1}(\Omega))\cap L^{\infty}(0,+\infty;(H^{1}(\Omega))^{\prime}),$
@@
$\mu\in L^{\infty}(0,+\infty;H^{1}(\Omega))\cap L_{\mathrm{uloc}}^{2}([0,+\infty);H^{3}(\Omega)\cap H^{2}_{N}(\Omega)),$
$\sigma\in L^{\infty}(0,+\infty;L^{6}(\Omega))\cap L_{\mathrm{uloc}}^{2}([0,+\infty);H^{1}(\Omega))\cap H^{1}(0,+\infty;(H^{1}(\Omega))^{\prime}).$
@@
$\partial_{t}\varphi+\bm{v}\cdot\nabla\varphi=\Delta\mu-\alpha(\overline{\varphi}-c_{0}),$
$\mu=-\Delta\varphi+\varPsi^{\prime}(\varphi)-\chi\sigma+\beta\mathcal{N}(\varphi-\overline{\varphi}),$
@@
$\left\langle\partial_{t}\sigma,\xi\right\rangle_{(H^{1})^{\prime},H^{1}}+(\bm{v}\cdot\nabla\sigma,\xi)+(\nabla\sigma,\nabla\xi)=\chi(\nabla\varphi,\nabla\xi),$
for almost all $t\in(0,+\infty)$ and all $\xi\in H^{1}(\Omega)$. The initial condition (3.3) is satisfied almost everywhere in $\Omega$. In addition, the solution $(\varphi,\mu,\sigma)$ satisfies the energy equality
@@
$\quad=\mathcal{F}(\varphi(t_{1}),\sigma(t_{1}))-\alpha\int_{t_{1}}^{t_{2}}\!\!\int_{\Omega}(\overline{\varphi}-c_{0})\mu\,\mathrm{d}x\mathrm{d}s+\int_{t_{1}}^{t_{2}}\!\!\int_{\Omega}\big[\bm{v}\cdot\nabla(\mu+\chi\sigma)\big]\varphi\,\mathrm{d}x\mathrm{d}s,$ (3.7)
for all $0\leq t_{1}\leq t_{2}<+\infty$, where $\mathcal{F}(\varphi,\sigma)$ is given by (1.8).
@@
For the single Cahn–Hilliard equation with non-smooth potential and divergence-free drift, existence and uniqueness of global weak solutions have been shown in [1, Theorem 6] and regularity properties under the mild requirement on the velocity have been established in [3, Theorem 2.4]. Theorem 3.1 and Theorem 3.2 generalize those previous results to the Cahn–Hilliard–diffusion system (3.1) with a cross-diffusion structure and a nonlocal interaction of Oono’s type. It is worth mentioning that under the condition $\bm{v}\in L^{\infty}(0,+\infty;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))\cap L^{2}(0,+\infty;\bm{H}_{0,\mathrm{div}}^{1}(\Omega))$ we can still gain full regularity of $(\varphi,\mu)$, but there is a loss in regularity of $\sigma$.